# Discovering Preference Optimization Algorithms with and for Large Language Models

### v2

@@ 6.1 Log Ratio Modulated Loss (DiscoPOP)
If the difference of log ratios is zero ($\rho=0)$, which is at the start of the training when the model policy $\pi_{\theta}$ is equal to the reference policy $\pi_{\text{ref}}$, then the loss is equally balanced between the logistic and exponential loss. If $\rho\rightarrow\infty$, the model policy diverges from the reference policy and chosen outputs are preferred, then the exponential term dominates. This emphasizes larger differences more strongly. On the other hand, if $\rho\rightarrow-\infty$, the model policy diverges from the reference policy and rejected outputs are preferred. In this case, the logistic loss can handle moderate differences well. The baseline objective losses and the LRML, the PADLL, and the AQFL functions are displayed in Figure 6, including their gradients. Surprisingly, we see that the DiscoPOP function has a non-convex segment and negative gradients at the starting point $\rho=0$. This is potentially helpful for introducing a curriculum or for stochasticity.
@@ 6.2 Limitations of DiscoPOP
In Figure 9 and Figure 10 of the Appendix, we plot the LRML objective function for $\beta\in\{0.01,0.025,0.05,0.1,0.25,0.5,1,2.5,5\}$ against DPO. When $\beta$ is high, the DiscoPOP objective function takes the form of the DPO log sigmoid loss. During training on $\beta=0.01$, we observed that DiscoPOP gets stuck in generating predominantly negative reviews (resulting in a reward score of $\sim 0.15$). We hypothesize that the loss is stuck in the local minima to the left with a negative difference in log ratios. While training with $\beta\in\{2.5,5.0\}$ we observed that the model collapsed after a sharp spike in the loss and subsequently having loss value 0 and NaN outputs. This is potentially due to a large gradient in the non-convex part, which could be amended with gradient clipping.
@@ 7 Related Work
Evolution and Search with Large Language Models. LLMs provide a fast and automated way to create multiple candidate solutions for a problem stated in natural language (Song et al., 2024), which makes them powerful tools for driving population-based search procedures. Various recent works have applied this approach to coding problems (Romera-Paredes et al., 2024), neural architecture search (Chen et al., 2024a), virtual robotic design settings (Lehman et al., 2023), reward functions (Ma et al., 2023; Yu et al., 2023), and algorithm heuristics (Liu et al., 2024). Finally, recently LLMs have shown to be capable of acting as recombination operators for black-box optimization with Evolution Strategies (Lange et al., 2024) and for Quality-Diversity approaches (Lim et al., 2024).
Automated Discovery for Machine Learning. There are many other approaches to automating the discovery of generalizable machine learning algorithms. Some prior works explore the space of ML functions using genetic algorithms and a hand-crafted domain-specific language for reinforcement learning algorithms (Co-Reyes et al., 2021), curiosity algorithms (Alet et al., 2020), and optimizers (Chen et al., 2024b). Other works instead parameterize a transferrable objective function using neural networks and optimize them with evolution strategies or meta-gradients. For example, Lu et al. (2022); Jackson et al. (2024b); Houthooft et al. (2018); Alfano et al. (2024); Kirsch et al. (2019); Oh et al. (2020); Jackson et al. (2024a) discover policy optimization objectives, Metz et al. (2022) evolves neural network optimizers, and Lange et al. (2023b, a) evolve blackbox optimizers.
@@ 8 Conclusion
Summary. In this paper, we proposed and used LLM-driven objective discovery to generate novel offline preference optimization algorithms. Specifically, we were able to discover high-performing preference optimization losses that achieve strong performance across held-out evaluation tasks, with the highest performing providing new insights into what an optimal objective may need to possess, such as being a blend of logistic and exponential losses, and possibly being non-convex.
Limitations & Future work. There are multiple limitations to our current approach. First, we have only scratched the surface of how to generate LLM objective proposals most effectively. Initial exploratory experiments using techniques such as temperature sampling or worst-to-best performance sorting in the context did not yield significant improvements. But one could imagine leveraging more information about the training runs and automatically tuning instruction prompt templates. E.g. by providing entire learning curve plots to a Visual Language Model (see Figure 12) or by meta-meta-optimizing (Lu et al., 2023) the LLM prompt. Second, the highest-performing loss re-purposed $\beta$ in the traditional sense, making it affect the functional behavior as well as the KL penalty of the model with respect to the base model. This motivates future work to study different forms, with perhaps multiple floating point parameters in the form, that each could be tuned separately. Although we provided an initial analysis sweep over this one single parameter and observed some instances of the functional behavior leading to instability of training the model, a further multi-parameter analysis, reformulating the objective, would be beneficial for future work. Finally, our work uses closed-source models (GPT-4) to generate code, which limits reproducibility and is costly to run. Future work could use the produced models themselves to generate code, resulting in code-level self-improvement.
@@
This work was supported by Azure sponsorship credits granted by Microsoft’s AI for Good Research Lab and by Microsoft’s Accelerate Foundation Models Academic Research initiative. The hardware used for training was sponsored by GoodAI. SH is funded by AstraZeneca. CF is funded by Canon Medical. AJC is funded by a Microsoft Research and EPSRC ICASE scholarship award. The code can also be accessed at https://github.com/samholt/DiscoPOP.
@@ B.1 Discovery Task - Single-turn Dialogue
For each valid generated objective function $f_{i}$, we use it to train an LLM and then collect a performance evaluation score. Specifically, we follow the same process when training and evaluating all objective functions, starting with a pre-trained supervised fine-tuned (SFT) 7 billion gemma model of ‘zephyr-7b-gemma-sft’ This is a 7 billion base version gemma [Gemma-Team et al., 2024] model supervised-fine-tuned on the ‘deita-10k-v0-sft’ dataset [Liu et al., 2023]. Starting with this model, we train it on the pairwise preference dataset of ‘Argilla DPO Mix 7K’; which attempts to create a high-quality preference dataset by filtering only highly rated chosen responses from the datasets of a multi-turn dataset, instruction following dataset [Longpre et al., 2023] and a diverse preference dataset that covers truthfulness, honesty and helpfulness [Cui et al., 2023]. For each training run, we trained all the parameters of the starting model, using a fixed $\beta=0.05$. We used the same fixed hyper-parameters for all training runs unless explicitly noted. Specifically, we used a learning rate of 5e-7, bfloat16 floating-point format, two epochs, a batch size per device of two, a gradient accumulation step of 8, a cosine learning rate scheduler, and AdamW optimization algorithm [Loshchilov and Hutter, 2017]. We use the popular TRL transformers library [von Werra et al., ], adapting the offline preference optimization objective function to train all models. The models were trained on 8 Nvidia A100 GPUs. An individual training run takes approximately 30 minutes. We provide training and evaluation statistics for discovered objective functions in Figure 7.
@@ C.1 MT-Bench
To assess the fitness of the discovered preference optimization loss function during the discovery phase, we evaluate the trained LLMs on the MT-Bench [Zheng et al., 2024] benchmark. The evaluation benchmark consists of high quality 80 multi-turn questions, from various disciplines. The goal is to assess LLM’s ability to follow instructions and keep the flow of a conversation. A larger LLM, in our case GPT-4, is then used as judge to score the quality of the answers with a number from 0 (lowest) to 10 (highest). Scores are given based on the quality of the LLM’s first-turn answer (single-turn), as well as on first and second answers (multi-turn). Finally, the MT-Bench score is the average of single-turn and multi-turn scores. For answer generation and evaluation, we used the FastChat library [^9] https://github.com/lm-sys/FastChat and its standard sampling and temperature parameters, provided by Zheng et al. [2024].
@@ C.2 Alpaca Eval
Currently, Alpaca Eval 2.0 [Li et al., 2023, Dubois et al., 2023, 2024] is also a popular benchmark to evaluate LLMs. This is a single-turn dialogue LLM-based automatic evaluation using a stronger LLM, here GPT-4 Turbo, to assess the win rate of the trained LLM policy’s completion compared to either GPT-4 or the of the underlying SFT base model. Specifically, Alpaca Eval 2.0, has been validated against 20K human annotations, and aims to reduce the length bias of Alpaca Eval; where using length-controlled (LC) Alpaca Eval shows a correlation with Chatbot Arena of 0.98, making it a popular benchmark with the highest correlation to Chatbot Arena [Dubois et al., 2024]. The Alpaca evaluation dataset consists of 841 high-quality instructions, originating from different data sets. The library [^10] https://github.com/tatsu-lab/alpaca_eval provided by Dubois et al. [2024] calculates the win-rate (percentage were the trained policy is prefered over the reference policy, first introduced in Alpaca Eval 1.0), and a length-controlled win-rate, where a linear model is fitted to de-bias for length of the prompt and instruction difficulty. To generate the answers we use a temperature of 0.7, sampling, and a maximum number of new tokens of 1024. Furthermore, the library provides the standard error of the mean, which indicates the confidence of the win-rate and LC win-rate.
@@ C.4 IMDb Rewards vs KL-Divergence
For the positive text generation, we do not require an LLM judge compared to MT-Bench, Alpaca Eval 2.0, and TL;DR evaluation, as we take a pre-trained sentiment classifier [^11] https://huggingface.co/siebert/sentiment-roberta-large-english as ground truth reward scorer. For the positive text generation, the LLMs apply sampling and a maximum of 60 new tokens. The rewards and KL-divergence are averaged over 10 different generations from the trained LLMs.
@@ D.3 Discovery Robustness with respect to LLM Hyperparameters
D.4 Visual Language Models for Objective Discovery
@@ Appendix E Discovered Objective Functions
Moreover, we display the code of the discovered losses as it is output by the LLM. In addition, we provide a mathematical representation of each, which we have adapted to be consistent with $\beta$ being the KL-Divergence regularization parameter. This is due to the fact, that the generated code for LRML, DBAQL, AQL, AQFL, and PFL did not uphold the $\beta$ ought to be multiplied with the difference of log-ratios, before any further calculations. If this was not upheld, it could to the loss function changing shapes based on the KL-regularization term, and therefore models could not converge, or potentially collapse. In future work, we should constrain the exploring LLM to uphold the $\beta$ multiplication with the input, before any other calculations are done with the difference of log-ratios $\rho$. As the meta exploration was done with a set $\beta=0.05$, and we wish to keep consistent with this scale of regularization, we have adapted the losses by dividing $\rho$ values used in intermediate calculations with a scalar $\tau=0.05$.
In the IMDb experiment in Section 5, we have thus used the corrected version of codes for the discovered losses, based on the provided mathematical representation, as we were most interested in the effect of the KL-divergence compared to the model rewards.
@@ E.7 PFL: Policy Focused Loss
Appendix F Full Run Log

### v3

@@ 6.1 Log Ratio Modulated Loss (DiscoPOP)
If the difference of log ratios is zero ($\rho=0)$, which is at the start of the training when the model policy $\pi_{\theta}$ is equal to the reference policy $\pi_{\text{ref}}$, then the loss is equally balanced between the logistic and exponential loss. If $\rho\rightarrow\infty$, the model policy diverges from the reference policy and chosen outputs are preferred, then the exponential term dominates. This emphasizes larger differences more strongly. On the other hand, if $\rho\rightarrow-\infty$, the model policy diverges from the reference policy, and rejected outputs are preferred. In this case, the logistic loss can handle moderate differences well. The baseline objective losses and the LRML, the PADLL, and the AQFL functions are displayed in Figure 6, including their gradients. Surprisingly, the DiscoPOP function has a non-convex segment and negative gradients at the starting point $\rho=0$. This is potentially helpful for introducing a curriculum or for stochasticity. Additional results and analysis of the discovered loss function can be found in Appendix D.6.
@@ 6.2 Limitations of DiscoPOP
In Figure 9 and Figure 10 of the Appendix, we plot the LRML objective function for $\beta\in\{0.01,0.025,0.05,0.1,0.25,0.5,1,2.5,5\}$ against DPO. When $\beta$ is high, the DiscoPOP objective function takes the form of the DPO log sigmoid loss. During training on $\beta=0.01$, we observed that DiscoPOP gets stuck in generating predominantly negative reviews (resulting in a reward score of $\sim 0.15$). We hypothesize that the loss is stuck in the local minima to the left with a negative difference in log ratios. While training with $\beta\in\{2.5,5.0\}$, we observed that the model collapsed after a sharp spike in the loss and subsequently having loss value 0 and NaN outputs. This is potentially due to a large gradient in the non-convex part, which could be amended with gradient clipping.
@@ 7 Related Work
Evolution and Search with Large Language Models. LLMs provide a fast and automated way to create multiple candidate solutions for a problem stated in natural language (Song et al., 2024), which makes them powerful tools for driving population-based search procedures. Various recent works have applied this approach to coding problems (Romera-Paredes et al., 2024), neural architecture search (Chen et al., 2024a; Holt et al., 2024b), virtual robotic design settings (Lehman et al., 2023), reward functions (Ma et al., 2023; Yu et al., 2023), and algorithm heuristics (Liu et al., 2024). Finally, recently LLMs have shown to be capable of acting as recombination operators for black-box optimization with Evolution Strategies (Lange et al., 2024) and for Quality-Diversity approaches (Lim et al., 2024).
Automated Discovery for Machine Learning. There are many other approaches to automating the discovery of generalizable machine learning algorithms. Some prior works explore the space of ML functions using genetic algorithms and a hand-crafted domain-specific language for reinforcement learning algorithms (Co-Reyes et al., 2021), curiosity algorithms (Alet et al., 2020), and optimizers (Chen et al., 2024b). Other works instead parameterize a transferrable objective function using neural networks and optimize them with evolution strategies or meta-gradients. For example, Lu et al. (2022); Jackson et al. (2024b); Houthooft et al. (2018); Alfano et al. (2024); Kirsch et al. (2019); Oh et al. (2020); Jackson et al. (2024a) discover policy optimization objectives, Metz et al. (2022) evolves neural network optimizers, and Lange et al. (2023b, a) evolve blackbox optimizers. Moreover, automatically discovering closed-form functions (i.e., symbolic regression), works exist that leverage RL (Petersen et al., 2020), gradient descent (Kacprzyk et al., 2024), RL with evolution strategies (Mundhenk et al., 2021), pre-training transformers (Biggio et al., 2021) and hybrid combinations of pre-training transformers, which are further refined with RL and evolution strategies (Holt et al., 2023b).
@@ 8 Conclusion
Summary. In this paper, we proposed and used LLM-driven objective discovery to generate novel offline preference optimization algorithms. Specifically, we were able to discover high-performing preference optimization losses that achieve strong performance across held-out evaluation tasks, with the highest performing providing new insights into what an optimal objective may need to possess, such as being a blend of logistic and exponential losses and possibly being non-convex.
Limitations & Future work. There are multiple limitations to our current approach. First, we have only scratched the surface of how to generate LLM objective proposals effectively. Initial exploratory experiments using techniques such as temperature sampling or worst-to-best performance sorting in the context did not yield significant improvements. But one could imagine leveraging more information about the training runs and automatically tuning instruction prompt templates. E.g. by providing entire learning curve plots to a Visual Language Model (see Figure 13) or by meta-meta-optimizing (Lu et al., 2023) the LLM prompt. Second, the highest-performing loss re-purposed $\beta$ in the traditional sense, making it affect the functional behavior and the KL penalty of the model with respect to the base model. This motivates future work to study different forms, with perhaps multiple floating point parameters in the form, that each could be tuned separately. Although we provided an initial analysis sweep over this one single parameter and observed some instances of the functional behavior leading to instability of training the model, a further multi-parameter analysis, reformulating the objective, would be beneficial for future work. Finally, our work uses closed-source models (GPT-4) to generate code, which limits reproducibility and is costly to run. Future work could use the produced models themselves to generate code, resulting in code-level self-improvement.
@@
This work was supported by Azure sponsorship credits granted by Microsoft’s AI for Good Research Lab and by Microsoft’s Accelerate Foundation Models Academic Research initiative. The hardware used for training was sponsored by GoodAI. SH is funded by AstraZeneca. AJC is funded by a Microsoft Research and EPSRC ICASE scholarship award. CL and RTL were supported by Sakana AI at the time of this work. The code can also be accessed at https://github.com/samholt/DiscoPOP.
@@ B.1 Discovery Task - Single-turn Dialogue
For each valid generated objective function $f_{i}$, we use it to train an LLM and then collect a performance evaluation score. Specifically, we follow the same process when training and evaluating all objective functions, starting with a pre-trained supervised fine-tuned (SFT) 7 billion gemma model of ‘zephyr-7b-gemma-sft’ This is a 7 billion base version gemma [Gemma-Team et al., 2024] model supervised-fine-tuned on the ‘deita-10k-v0-sft’ dataset [Liu et al., 2023]. Starting with this model, we train it on the pairwise preference dataset of ‘Argilla DPO Mix 7K’; which attempts to create a high-quality preference dataset by filtering only highly rated chosen responses from the datasets of a multi-turn dataset, instruction following dataset [Longpre et al., 2023] and a diverse preference dataset that covers truthfulness, honesty and helpfulness [Cui et al., 2023]. For each training run, we trained all the parameters of the starting model, using a fixed $\beta=0.05$. We used the same fixed hyper-parameters for all training runs unless explicitly noted. Specifically, we used a learning rate of 5e-7, bfloat16 floating-point format, two epochs, a batch size per device of two, a gradient accumulation step of 8, a cosine learning rate scheduler, and AdamW optimization algorithm [Loshchilov and Hutter, 2017]. We use the popular TRL transformers library [von Werra et al., ], adapting the offline preference optimization objective function to train all models. The models were trained on 8 Nvidia A100 GPUs. An individual training run takes approximately 30 minutes. We provide training and evaluation statistics for discovered objective functions in Figure 7. We also provide the equivalent code implementation at https://github.com/vanderschaarlab/DiscoPOP.
@@ C.1 MT-Bench
To assess the fitness of the discovered preference optimization loss function during the discovery phase, we evaluate the trained LLMs on the MT-Bench [Zheng et al., 2024] benchmark. The evaluation benchmark consists of 80 high-quality multi-turn questions from various disciplines. The goal is to assess LLM’s ability to follow instructions and keep the flow of a conversation. A larger LLM, in our case GPT-4, is then used as a judge to score the quality of the answers with a number from 0 (lowest) to 10 (highest). Scores are given based on the quality of the LLM’s first-turn answer (single-turn) and first and second answers (multi-turn). Finally, the MT-Bench score is the average of single-turn and multi-turn scores. For answer generation and evaluation, we used the FastChat library [^9] https://github.com/lm-sys/FastChat and its standard sampling and temperature parameters, provided by Zheng et al. [2024].
@@ C.2 Alpaca Eval
Alpaca Eval 2.0 [Li et al., 2023, Dubois et al., 2023, 2024] is also a popular benchmark for evaluating LLMs. This is a single-turn dialogue LLM-based automatic evaluation using a stronger LLM, here GPT-4 Turbo, to assess the win rate of the trained LLM policy’s completion compared to either GPT-4 or the of the underlying SFT base model. Specifically, Alpaca Eval 2.0 has been validated against 20K human annotations and aims to reduce the length bias of Alpaca Eval, where using length-controlled (LC) Alpaca Eval shows a correlation with Chatbot Arena of 0.98, making it a popular benchmark with the highest correlation to Chatbot Arena [Dubois et al., 2024]. The Alpaca evaluation dataset consists of 841 high-quality instructions from different data sets. The library [^10] https://github.com/tatsu-lab/alpaca_eval provided by Dubois et al. [2024] calculates the win-rate (percentage were the trained policy is prefered over the reference policy, first introduced in Alpaca Eval 1.0), and a length-controlled win-rate, where a linear model is fitted to de-bias for length of the prompt and instruction difficulty. We used a temperature of 0.7, sampling, and a maximum number of new tokens of 1024 to generate the answers. Furthermore, the library provides the standard error of the mean, which indicates the confidence of the win-rate and LC win-rate.
@@ C.4 IMDb Rewards vs KL-Divergence
For the positive text generation, we do not require an LLM judge compared to MT-Bench, Alpaca Eval 2.0, and TL;DR evaluation, as we take a pre-trained sentiment classifier [^11] https://huggingface.co/siebert/sentiment-roberta-large-english as ground truth reward scorer. The LLMs apply sampling and a maximum of 60 new tokens for the positive text generation. The rewards and KL-divergence are averaged over 10 different generations from the trained LLMs.
@@ D.3 Discovery Robustness with respect to LLM Hyperparameters
D.4 LLM-Driven Discovery Analysis
Figure 12: LLM-driven discovery for CIFAR-10 loss functions with (left) and without (right) providing fitness feedback. Note that without feedback, it performs worse but also is unable to refine its ideas, resulting in fewer improvements across generations.
D.5 Visual Language Models for Objective Discovery
@@
D.6 Additional Analysis of DiscoPOP
We performed further analysis to improve our understanding of the discovered loss function. We hypothesize that the local optimum of the DiscoPOP loss could catch noisy or incorrect data points. By inspecting the DiscoPOP loss values and log ratios of the training preference pairs in the IMDb dataset, we see that 1.35% of training points fall there (see Figure 14). Although we use the binary preference labels from the IMDb dataset [^12] https://huggingface.co/datasets/ZHZisZZ/imdb_preference for training, the dataset also includes a positivity reward score for each completion, calculated by a separate reward model. When we analyze the data points between the local optima, we find that the positive and negative completions are significantly closer in absolute reward difference than those outside the local optimum (See Table 4). This implies that the preference labels on those points are likely more challenging to distinguish and help empirically validate our hypothesis.
Figure 14: Distribution of $\beta$-scaled difference of log-ratios (left y-axis) and corresponding DiscoPOP loss value (right y-axis) of the training samples on the IMDb positive review generation task. The DiscoPOP function has a local minimum at $f_{lrml}(-2.3714)=0.785929$ and a local maximum at $f_{lrml}(1.44012)=0.87829$. The number of samples within the two local optima corresponds to 1.35% of the training set.
Description Domain $\beta\rho$ Mean 95%-CI $p$-value % of training set
Between optima $[-2.3714,1.44012]$ 0.981 (0.830, 1.131) - 1.35
Full range $\mathbb{R}$ 3.861 (3.818, 3.904) $<0.01$% 100
Outside optima \$\mathbb{R}\backslash[-2.3714,1.44012]$ 3.9 (3.857, 3.944) $<0.01$% 98.65
Below local minimum $(-\infty,-2.3714)$ 4.086 (4.041, 4.131) $<0.01$% 6.22
Above local maximum $(1.44012,\infty)$ 1.141 (1.076, 1.206) 4.29% 92.43
Table 4: The IMDb positive review preference dataset also provides “golden reward” scores for the chosen responses $r_{w}$ and rejected responses $r_{l}$, calculated with a strong sentiment classifier ($\log p(\text{pos})-\log p(\text{neg})$). We analysed the absolute difference in reward scores between the chosen and rejected responses $|r_{w}-r_{l}|$ across the different domains of the DiscoPOP function and report the statistics here. Training samples within the optima have a significantly lower mean absolute difference in rewards compared to the other regions in the DiscoPOP loss function. The samples stuck within the optima are the training triplets where the chosen and rejected responses are “closest” to each other and likely the most “noisy” labels.
@@ Appendix E Discovered Objective Functions
Moreover, we display the code of the discovered losses as it is output by the LLM. In addition, we provide a mathematical representation of each, which we have adapted to be consistent with $\beta$ being the KL-Divergence regularization parameter. This is because the generated code for LRML, DBAQL, AQL, AQFL, and PFL did not uphold the $\beta$ ought to be multiplied with the difference of log-ratios before any further calculations. If this was not upheld, it could lead to the loss function changing shapes based on the KL-regularization term, and therefore, models could not converge or potentially collapse. In future work, we should constrain the exploring LLM to uphold the $\beta$ multiplication with the input before any other calculations are done with the difference of log-ratios $\rho$. As the meta-exploration was done with a set $\beta=0.05$, and we wish to keep consistent with this scale of regularization, we have adapted the losses by dividing $\rho$ values used in intermediate calculations with a scalar $\tau=0.05$.
In the IMDb experiment in Section 5, we have thus used the corrected version of codes for the discovered losses based on the provided mathematical representation, as we were most interested in the effect of the KL-divergence compared to the model rewards.
@@ E.7 PFL: Policy Focused Loss
Appendix F Extended Related Work
In the following, we provide an extended related work of the related work included in the main paper.
Evolution and Search with Large Language Models. LLMs provide a fast and automated way to create multiple candidate solutions for a problem stated in natural language [Song et al., 2024], which makes them powerful tools for driving population-based search procedures. Various recent works have applied this approach to coding problems [Romera-Paredes et al., 2024], neural architecture search [Chen et al., 2024a, Holt et al., 2024b, d], virtual robotic design settings [Lehman et al., 2023], reward functions [Ma et al., 2023, Yu et al., 2023], and algorithm heuristics [Liu et al., 2024]. Finally, recently LLMs have shown to be capable of acting as recombination operators for black-box optimization with Evolution Strategies [Lange et al., 2024] and for Quality-Diversity approaches [Lim et al., 2024]. Additionally, prior work has shown LLM multi-agent frameworks’ ability to excel at large code-generation tasks [Holt et al., 2024c], which we could envision enabling scaling up searching for larger candidate solutions for more complicated objectives or algorithms than the existing approaches that fit within the LLMs output context window.
Automated Discovery for Machine Learning. There are many other approaches to automating the discovery of generalizable machine learning algorithms. Some prior works explore the space of ML functions using genetic algorithms and a hand-crafted domain-specific language for reinforcement learning algorithms [Co-Reyes et al., 2021], curiosity algorithms [Alet et al., 2020], and optimizers [Chen et al., 2024b]. Other works instead parameterize a transferrable objective function using neural networks and optimize them with evolution strategies or meta-gradients. For example, Lu et al. [2022], Jackson et al. [2024b], Houthooft et al. [2018], Alfano et al. [2024], Kirsch et al. [2019], Oh et al. [2020], Jackson et al. [2024a] discover policy optimization objectives, Metz et al. [2022] evolves neural network optimizers, and Lange et al. [2023b, a] evolve blackbox optimizers. Moreover, automatically discovering closed-form functions (i.e., symbolic regression), works exist that leverage RL [Petersen et al., 2020], gradient descent [Kacprzyk et al., 2024], RL with evolution strategies [Mundhenk et al., 2021], pre-training transformers [Biggio et al., 2021] and hybrid combinations of pre-training transformers, which are further refined with RL and evolution strategies [Holt et al., 2023b]. Furthermore, more general-purpose ML approaches exist to perform sequential modeling of processes [Holt et al., 2022], which can be readily used in model-based RL approaches [Holt et al., 2023a, 2024a].
Preference Optimization Algorithms. While the reduction to supervised learning makes DPO and alternatives easier to use, other approaches have sought to simplify the RL step, including using variants of REINFORCE [Ahmadian et al., 2024, Gemma-Team et al., 2024] as well as more fine-grained feedback [Wu et al., 2024] through preferences over individual steps in the reasoning process [Uesato et al., 2022, Lightman et al., 2023] or reward redistribution [Chan et al., 2024]. Others use iterative offline training interleaved with sampling from the policy model and obtaining a preference ranking from themselves [Xu et al., 2023], another judge LLM [Guo et al., 2024], or an oracle [Swamy et al., 2024].
Appendix G Full Run Log